### v1

# Information Propagation in Predator-Prey Dynamics of Turbulent Plasma

Abstract
Magnetically confined fusion plasmas exhibit predator-prey-like cyclic oscillations through the self-regulating interaction between drift-wave turbulence and zonal flow. To elucidate the detailed mechanism and causality underlying this phenomenon, we construct a simple stochastic predator-prey model that incorporates intrinsic fluctuations and analyze its statistical properties from an information-theoretic perspective. We first show that the model exhibits persistent fluctuating cyclic oscillations called quasi-cycles due to amplification of intrinsic noise. This result suggests the possibility that the previously observed periodic oscillations in a toroidal plasma are not limit cycles but quasi-cycles, and that such quasi-cycles may be widely observed under various conditions. For this model, we further prove that information of zonal flow is propagated to turbulence. This information-theoretic analysis may provide a theoretical basis for regulating turbulence by controlling zonal flow.
pacs:
Valid PACS appear here
Introduction.
—Predator-prey dynamics can be observed ubiquitously in various systems such as ecosystems [1, 2, 3], genetic systems [4], fluids [5, 6, 7, 8], or even quantum systems [9]. Magnetically confined fusion plasmas also exhibit predator-prey-like cyclic oscillations in various situations, including the self-regulating process between zonal flow and drift-wave turbulence [10, 11, 12, 13], the intermediate phase (I-phase) during the low-to-high confinement (L-H) transition [14, 15, 16, 17, 18], the formation and collapse of internal transport barriers [19], and self-regulated oscillations of magnetic islands [20]. Among these, the interplay between zonal flow (predator) and drift-wave turbulence (prey) is believed to play a crucial role in suppressing anomalous heat and particle transport, and therefore, intensive attempts have been made to regulate turbulence by controlling zonal flow. Despite this importance, the detailed mechanism and causality underlying the cyclic oscillations have not been elucidated [12]. To deepen our understanding of the self-regulating process between zonal flow and turbulence, it is desirable to construct a minimal model that exhibits the cyclic oscillations and investigate the causality within the oscillations.
The main aim of this Letter is twofold. The first is to construct a simple model for the cyclic oscillations observed in plasma turbulence. Note that the standard two-variable predator-prey model proposed by Diamond et al. [21, 10], which incorporates the self-regulation mechanism between turbulence and zonal flow, does not exhibit a stable cyclic oscillation. To address this issue, we aim to construct a stochastic predator-prey model that incorporates the effects of intrinsic noise from the perspective of statistical physics. Since it has been pointed out in the field of ecology that it is essential to consider intrinsic fluctuations in predator-prey dynamics [22, 23], we expect that such a statistical physics approach provides useful insights into plasma turbulence. The second is to quantify the statistical causality between zonal flow and turbulence in this model through the lens of information theory. While several previous studies have attempted to quantify the statistical causality in plasma turbulence using information-theoretic quantities, such as (net) transfer entropy [24, 25] or information length [26, 27, 28, 29, 30, 31, 32], these quantities do not change sign under time reversal and thus can be nonzero even in an equilibrium state where all probability currents vanish. Instead, here we employ information flow, which is a pivotal concept within the framework of information thermodynamics [33, 34, 35, 36, 35, 37, 38] and has recently been applied to standard fluid turbulence [39, 40, 41].
The constructed model can be interpreted as a standard two-variable predator-prey model that takes into account intrinsic fluctuations. We show that the model exhibits persistent fluctuating cyclic oscillations, called quasi-cycles [22, 23], due to stochastic amplification. The quasi-cycles can be observed for a wide range of parameters, including parameter values consistent with experiments [14] and gyrokinetic simulations [42]. This result suggests the possibility that the previously observed periodic oscillations in a toroidal plasma are not limit cycles, but rather quasi-cycles. Our analysis of the resonance condition also implies that such quasi-cycles may be widely observed under various conditions.
For this model, we quantify the causality between zonal flow and turbulence using information flow. Specifically, we analytically show that the information on zonal flow is propagated to turbulence. This information propagation occurs as long as zonal flow and turbulence coexist, regardless of the presence or absence of quasi-cycles. Thus, our information-theoretic analysis may provide a theoretical foundation for the regulation of turbulence through the control of zonal flow.
Figure 1: (a) Schematic of the model. (b) Vector fields for parameter values $p=0.5,d=b=\gamma=1$ (left) and $p=2,d=b=\gamma=1$ (right). The color bar denotes the magnitude of the vector fields. The red markers indicate the stable fixed points, and the blue markers indicate the unstable fixed points. (c) The trajectory of the Langevin equation. The parameter values are $\epsilon=1/2,\beta=1,D^{W}=1/2$. The characteristic period is given by $2\pi/\omega_{c}\simeq 10$.
Setup.
[…]
$\partial_{t}p_{t}({\bm{n}})$ $=\sum_{\rho}\left[W_{\rho}({\bm{n}}|{\bm{n}}-{\bm{S}}_{\rho})p_{t}({\bm{n}}-{\bm{S}}_{\rho})\right$.
$\qquad\qquad\qquad\left.-W_{\rho}({\bm{n}}+{\bm{S}}_{\rho}|{\bm{n}})p_{t}({\bm{n}})\right].$ (1)
Here, $W_{\rho}({\bm{n}}+{\bm{S}}_{\rho}|{\bm{n}})$ denotes the transition rate from state $\bm{n}$} to state ${\bm{n}}+{\bm{S}}_{\rho}$ due to the reaction $\rho$, where ${\bm{S}}_{\rho}=(S^{Z}_{\rho},S^{W}_{\rho})$ denotes the stoichiometric vectors, which quantifies the change of $\bm{n}$} during a reaction of type $\rho$. For simplicity, we assume the mass action law [43]:
$W_{\rho}({\bm{n}}+{\bm{S}}_{\rho}|{\bm{n}})$ $=\Omega\rho\prod_{\alpha\in\{Z,W\}}\dfrac{1}{\Omega^{\nu^{\alpha}_{\rho}}}\dfrac{n^{\alpha}!}{(n^{\alpha}-\nu^{\alpha}_{\rho})!},$ (2)
[…]
System size expansion.
—Since we are interested in the collective behavior of the system, we consider coarse-grained descriptions by taking the infinite volume limit $\Omega\rightarrow\infty$. Here, we note that the two limits $t\rightarrow\infty$ and $\Omega\rightarrow\infty$ do not commute in this model, which is known as Keizer’s paradox [44, 45]. Indeed, if we take the limit $t\rightarrow\infty$ first, then the probability distribution converges to a unique stationary distribution $p_{\mathrm{ss}}({\bm{n}})=\delta_{{\bm{n}},{\bm{0}}}$, which implies that eventually there will be no zonal flows and turbulences. Below, we focus on the opposite case, where we first take $\Omega\rightarrow\infty$ and then $t\rightarrow\infty$. The coarse-grained dynamics in the limit $\Omega\rightarrow\infty$ can be systematically obtained by applying van Kampen’s system size expansion [46]. To separate the fluctuating part from the average part, we introduce a new stochastic variable ${\bm{r}}_{t}=(z_{t},w_{t})\in\mathbb{R}^{2}$ as ${\bm{n}}_{t}/\Omega={\bm{c}}_{t}+{\bm{r}}_{t}/\Omega^{1/2}$, where ${\bm{c}}_{t}=(c^{Z}_{t},c^{W}_{t})\in\mathbb{R}^{2}_{\geq 0}$ denotes the mean energy density to be determined. By substituting this expression into Eq. (1) and expanding in terms of $\Omega^{-1/2}$, we find that the leading order $O(\Omega^{1/2})$ yields the time evolution equation for the average part ${\bm{c}}_{t}$, which can be regarded as a coarse-grained dynamics of Eq. (1):
$\dfrac{d}{dt}c^{Z}_{t}$ $=pc^{Z}_{t}c^{W}_{t}-dc^{Z}_{t},$ (3)
$\dfrac{d}{dt}c^{W}_{t}$ $=bc^{W}_{t}-\gamma(c^{W}_{t})^{2}-pc^{Z}_{t}c^{W}_{t}.$ (4)
See Ref. [47] for the detailed derivation. This equation corresponds to the standard two-variable predator-prey model proposed by Diamond et al. [21, 10], which has the same form as the Lotka–Volterra equation. There are two regimes in this model (Fig. 1(b)). When $bp<d\gamma$, the stable fixed point is $(c^{Z}_{\mathrm{ss}},c^{W}_{\mathrm{ss}})=(0,b/\gamma)$, which represents a state without zonal flows. In contrast, when $bp>d\gamma$, the stable fixed point is $((bp-d\gamma)/p^{2},d/p)$, which represents a coexistence state. Importantly, this equation does not exhibit limit cycles, contrary to the naive expectation of the existence of predator-prey oscillations. While cyclic oscillations emerge when $\gamma=0$, these cycles are not limit cycles but structurally unstable.
Note that the coarse-grained description [Eqs. (3) and (4)] ignores intrinsic noise that is inevitably present in the original individual-level description (1). For a realistic situation where $\Omega$ is large but finite, it becomes essential to consider the intrinsic noise. To investigate the effect of the intrinsic noise on the coexistence state, we focus on the regime $bp>d\gamma$ and proceed to the subleading order $O(\Omega^{0})$ of the system size expansion by substituting $(c^{Z}_{\mathrm{ss}},c^{W}_{\mathrm{ss}})=((bp-d\gamma)/p^{2},d/p)$. The resulting equation describes the time evolution of the fluctuating part ${\bm{r}}_{t}$ around ${\bm{c}}_{t}={\bm{c}}_{\mathrm{ss}}$ [47]:
$\dfrac{d}{d\tau}z_{\tau}$ $=\epsilon w_{\tau}+\zeta^{Z}_{\tau},$ (5)
$\dfrac{d}{d\tau}w_{\tau}$ $=-\beta z_{\tau}-(1-\epsilon)w_{\tau}+\zeta^{W}_{\tau},$ (6)
where $\tau:=bt$ denotes the dimensionless time, $\beta:=d/b$ denotes the dimensionless damping rate, and $\epsilon:=1-d\gamma/bp\in(0,1)$ denotes the ratio of time scales for $z$ and $w$. The terms $\zeta^{Z}_{\tau}$ and $\zeta^{W}_{\tau}$ denote the zero-mean white Gaussian noise that satisfies $\langle\zeta^{Z}_{\tau}\zeta^{Z}_{\tau^{\prime}}\rangle=2D^{Z}\delta(\tau-\tau^{\prime})$, $\langle\zeta^{Z}_{\tau}\zeta^{W}_{\tau^{\prime}}\rangle=-D^{Z}\delta(\tau-\tau^{\prime})$, and $\langle\zeta^{W}_{\tau}\zeta^{W}_{\tau^{\prime}}\rangle=2D^{W}\delta(\tau-\tau^{\prime})$, where $D^{Z}:=\epsilon c^{W}_{\mathrm{ss}}=\epsilon d/p$ and $D^{W}:=c^{W}_{\mathrm{ss}}=d/p$. Note that this linear Langevin equation has a form in which noise is added to the linearized Lotka–Volterra equation around the coexistence fixed point. We emphasize that the noise intensities $D^{Z}$ and $D^{W}$ are uniquely determined from the master equation (1). In other words, the noise driving the system is not artificially added but rather arises from intrinsic stochasticity. This property marks a stark difference from the stochastic models used in Refs. [26, 27, 28, 29, 30, 31, 32], where noise is artificially added to deterministic predator-prey models. The Langevin equations (5) and (6) are equivalent to the Fokker–Planck equation $\partial_{\tau}p_{\tau}({\bm{r}})=-\partial_{z}J^{Z}_{\tau}({\bm{r}})-\partial_{w}J^{W}_{\tau}({\bm{r}})$, where $p_{\tau}({\bm{r}})$ denotes the probability density of finding ${\bm{r}}_{\tau}={\bm{r}}$ and $J^{Z}_{\tau}({\bm{r}})$ and $J^{W}_{\tau}({\bm{r}})$ denote the probability currents [47].
Quasi-cycles induced by stochastic amplification.
—Although the coarse-grained deterministic description [Eqs.(3) and (4)] cannot predict stable cycles, the stochastic description [Eqs. (5) and (6)] predicts persistent cyclic oscillations called quasi-cycles. Figure 1(c) shows an example of such persistent fluctuating oscillations observed in the linear Langevin equation. Notably, the fluctuations of zonal flow follow those of turbulence with a phase lag $\sim\pi/2$, which are similar to the predator-prey-like oscillations observed in a toroidal plasma [14, 15, 16, 17, 12]. This oscillatory behavior occurs when the coexistence state is a stable spiral fixed point in the coarse-grained deterministic description. More precisely, from the calculation of the power spectral density [47], we can show that the resonant stochastic amplification occurs for a wide range of parameters, where the characteristic resonant frequency $\omega_{c}:=\sqrt{\epsilon\beta-(1-\epsilon)^{2}/2}$ takes a real value (see also Fig. 2(b)). Because the white noise in the Langevin equation covers all frequencies, the corresponding cycle with the resonant frequency $\omega_{c}$ is amplified. This mechanism is called stochastic amplification to distinguish it from stochastic resonance [22].
The quasi-cycles differ from limit cycles in that they are stochastic oscillations where the different frequencies near $\omega_{c}$ are all excited [23, 48]. Furthermore, because quasi-cycles can occur as long as the system has a stable spiral fixed point and noise, they may be widely observed in plasma turbulence under various conditions. Although we have emphasized that our model is a qualitative description, here we present a set of parameter values consistent with experiments and simulations for reference; By comparing with the gyrokinetic simulations [42], they can be estimated as $\epsilon\sim 1$ and $\beta\sim 10^{-4}$, which satisfies the resonance condition, and the corresponding cycle frequency reads $\omega_{c}\sim 10^{-2}$ in dimensionless units.
Information propagation between zonal flow and turbulence.
—The presence of quasi-cycles suggests that intrinsic fluctuations are essential for understanding the dynamics of zonal flow and turbulence. Therefore, we scrutinize the statistical causality underlying the intrinsic fluctuations from an information-theoretic viewpoint.
We first introduce the mutual information [49], which quantifies the mutual dependence between zonal flow and turbulence:
$I[Z\colon\!W]:=\int dzdwp_{\tau}(z,w)\ln\dfrac{p_{\tau}(z,w)}{p^{Z}_{\tau}(z)p^{W}_{\tau}(w)}\geq 0,$ (7)
where $p_{\tau}(z,w)$ denotes the joint probability density, and $p^{Z}_{\tau}(z)$ and $p^{W}_{\tau}(w)$ denote the marginal distributions. The mutual information is nonnegative and equal to zero if and only if zonal flow and turbulence are statistically independent.
Figure 2: (a) Schematic of information flow $\dot{I}^{Z\rightarrow W}_{\lambda}$. (b) $(\epsilon,\beta)$-dependence of $\dot{I}^{Z\rightarrow W}_{\lambda=1/2}$. The dashed line denotes the resonance condition $2\epsilon\beta-(1-\epsilon)^{2}=0$. The stochastic amplification occurs in the parameter region above this line.
The decomposition of the time derivative of the mutual information yields information flow. Specifically, the information flow from zonal flow to turbulence is defined as (see Fig. 2(a))
$\dot{I}^{Z\rightarrow W}_{\lambda}:=\lim_{\Delta\tau\rightarrow 0^{+}}\dfrac{I[Z_{\tau+\lambda\Delta\tau}\colon\!W_{\tau+\Delta\tau}]-I[Z_{\tau+\lambda\Delta\tau}\colon\!W_{\tau}]}{\Delta\tau}$ (8)
with $\lambda\in[0,1]$. We can similarly define the information flow from turbulence to zonal flow $\dot{I}^{W\rightarrow Z}_{\lambda}$, which is related to $\dot{I}^{Z\rightarrow W}_{\lambda}$ as $d_{\tau}I[Z\colon\!W]=\dot{I}^{Z\rightarrow W}_{\lambda}+\dot{I}^{W\rightarrow Z}_{1-\lambda}$. Hence, we have $\dot{I}^{Z\rightarrow W}_{\lambda}=-\dot{I}^{W\rightarrow Z}_{1-\lambda}$ in the steady state. Here, the information flow $\dot{I}^{Z\rightarrow W}_{\lambda}$ depends on $\lambda$ because the Langevin equations (5) and (6) do not satisfy the bipartite condition. That is, the noises in Eqs. (5) and (6) are not independent, but correlated as $\langle\zeta^{Z}_{\tau}\zeta^{W}_{\tau^{\prime}}\rangle=-D^{Z}\delta(\tau-\tau^{\prime})$. We remark that, in the context of information thermodynamics [35, 37, 38], it is common to consider bipartite systems, where information flow does not depend on $\lambda$, and non-bipartite systems have rarely been studied [36].
The information flow $\dot{I}^{Z\rightarrow W}_{\lambda}$ quantifies the rate at which turbulence gains information about zonal flow. In particular, its sign indicates the direction in which the information is transmitted: if $\dot{I}^{Z\rightarrow W}_{\lambda}>0$ for all $\lambda$, the information of zonal flow is transferred to turbulence, and vice versa. Here, we note that the non-bipartite structure allows the case where the sign of $\dot{I}^{Z\rightarrow W}_{\lambda}$ varies depending on $\lambda$, which makes it difficult to determine the direction of information transfer [36]. In that case, $\dot{I}^{Z\rightarrow W}_{\lambda=1/2}$ plays a special role because it is odd under time reversal and vanishes in the equilibrium state. Fortunately, as will be shown below, this problem is irrelevant to our model.
From Eq. (8), we can express the information flow in terms of the probability current $J^{W}_{\tau}({\bm{r}})$ as
$\dot{I}^{Z\rightarrow W}_{\lambda}$ $=\int dzdw\left[J^{W}_{\tau}({\bm{r}})+(2\lambda-1)\dfrac{D^{Z}}{2}\dfrac{\partial}{\partial z}p_{\tau}({\bm{r}})\right]$
$\qquad\times\dfrac{\partial}{\partial w}\ln\dfrac{p_{\tau}({\bm{r}})}{p_{\tau}(z)p_{\tau}(w)},$ (9)
where $p_{\tau}(z)$ and $p_{\tau}(w)$ are marginal distributions for zonal flow and turbulence, respectively. By noting that the steady-state distribution $p_{\mathrm{ss}}({\bm{r}})$ for the linear Langevin equations (5) and (6) is Gaussian with covariance matrix $\Sigma:=\langle({\bm{r}}-\langle{\bm{r}}\rangle)({\bm{r}}-\langle{\bm{r}}\rangle)^{\top}\rangle$, the information flow in the steady state can be calculated as
$\dot{I}^{Z\rightarrow W}_{\lambda}$ $=\dfrac{D^{Z}}{\det\Sigma}\left[\Sigma^{WW}+(1-\lambda)\Sigma^{ZW}\right]$
$>0\quad\text{for all}\quad\lambda\in[0,1],$ (10)
where $\Sigma^{ZW}=\langle(z-\langle z\rangle)(w-\langle w\rangle)\rangle$ and $\Sigma^{WW}=\langle(w-\langle w\rangle)^{2}\rangle$. See Ref. [47] for the detailed derivation of Eqs. (9) and (10). The inequality (10) states that the information about fluctuations of zonal flow is propagated to turbulence. In other words, the turbulence is “learning” about the fluctuating behavior of the zonal flow. This result is somewhat counterintuitive because the system exhibits persistent predator-prey cycles, where the fluctuations of zonal flow follow those of turbulence with a time lag. Figure 2(b) shows the $(\epsilon,\beta)$-dependence of $\dot{I}^{Z\rightarrow W}_{\lambda=1/2}$. Interestingly, Fig. 2(b) suggests that the behavior of the information flow does not depend significantly on whether stochastic amplification occurs or not. In other words, even when the system does not exhibit the fluctuating predator-prey oscillation, there is a statistical causality from zonal flow to turbulence. This information-theoretic analysis may provide a theoretical foundation for the regulation of turbulence through the control of zonal flow. We finally remark that the inferred causality from the information flow is consistent with that from the cross-correlation function, although interpreting the latter is less straightforward [47].
Concluding remarks.
—We have proposed a simple stochastic predator-prey model for plasma turbulence that exhibits quasi-cycles induced by stochastic amplification of intrinsic noise. For this model, we have quantified the causality between zonal flow and turbulence by showing that the information of zonal flow is propagated to turbulence.
Our model illustrates the significance of considering the effects of intrinsic noise in understanding the predator-prey-like oscillations observed in plasma turbulence. Notably, our results suggest the possibility that the observed cyclic oscillations in a toroidal plasma are quasi-cycles rather than limit cycles. Using the methods proposed in Ref. [48], it would be possible to distinguish quasi-cycles and limit cycles both in experiments and simulations. Further studies are needed to verify whether both quasi-cycles and information transfer from zonal flow to turbulence can be observed in experiments and simulations of more realistic models, such as the Hasegawa–Wakatani equation [50].
Acknowledgements.
We thank S. Maeyama and M. Nakata for fruitful discussions. T.T. was supported by JSPS KAKENHI Grant Number JP25K17315 and JST PRESTO Grant Number JPMJPR23O6, Japan. M.S. was supported by JSPS KAKENHI Grant Number JP25K00986, Japan.

### v2

@@
 This result suggests the possibility that the previously observed periodic oscillations in a toroidal plasma are not limit cycles but quasi-cycles, and that such quasi-cycles may be widely observed under various conditions.
-For this model, we further prove that information of zonal flow is propagated to turbulence.
-This information-theoretic analysis may provide a theoretical basis for regulating turbulence by controlling zonal flow.
+For this model, we further prove that information of the zonal flow is propagated to turbulence.
+This result suggests that turbulence behavior may be predictable to a certain extent based on zonal flow characteristics.
 pacs:
 Valid PACS appear here
-Introduction.
-—Predator-prey dynamics can be observed ubiquitously in various systems such as ecosystems [1, 2, 3], genetic systems [4], fluids [5, 6, 7, 8], or even quantum systems [9].
+I Introduction
+Predator-prey dynamics can be observed ubiquitously in various systems such as ecosystems [1, 2, 3], genetic systems [4], fluids [5, 6, 7, 8], or even quantum systems [9].
 Magnetically confined fusion plasmas also exhibit predator-prey-like cyclic oscillations in various situations, including the self-regulating process between zonal flow and drift-wave turbulence [10, 11, 12, 13], the intermediate phase (I-phase) during the low-to-high confinement (L-H) transition [14, 15, 16, 17, 18], the formation and collapse of internal transport barriers [19], and self-regulated oscillations of magnetic islands [20].
@@
 The main aim of this Letter is twofold.
-The first is to construct a simple model for the cyclic oscillations observed in plasma turbulence.
+The first is to construct a simple qualitative model for the cyclic oscillations observed in plasma turbulence.
 Note that the standard two-variable predator-prey model proposed by Diamond et al.
@@
 To address this issue, we aim to construct a stochastic predator-prey model that incorporates the effects of intrinsic noise from the perspective of statistical physics.
-Since it has been pointed out in the field of ecology that it is essential to consider intrinsic fluctuations in predator-prey dynamics [22, 23], we expect that such a statistical physics approach provides useful insights into plasma turbulence.
-The second is to quantify the statistical causality between zonal flow and turbulence in this model through the lens of information theory.
-While several previous studies have attempted to quantify the statistical causality in plasma turbulence using information-theoretic quantities, such as (net) transfer entropy [24, 25] or information length [26, 27, 28, 29, 30, 31, 32], these quantities do not change sign under time reversal and thus can be nonzero even in an equilibrium state where all probability currents vanish.
-Instead, here we employ information flow, which is a pivotal concept within the framework of information thermodynamics [33, 34, 35, 36, 35, 37, 38] and has recently been applied to standard fluid turbulence [39, 40, 41].
+Since it has been pointed out in the field of ecology that it is essential to consider intrinsic fluctuations in predator-prey dynamics [22, 23], we expect that such a statistical physics approach provides valuable insights into plasma turbulence.
+The second is to quantify the predictive causality between zonal flow and turbulence in this model through the lens of information theory.
+Several previous studies have attempted to quantify the predictive causality in plasma turbulence using information-theoretic quantities, such as (net) transfer entropy [24, 25] or information length [26, 27, 28, 29, 30, 31, 32, 33].
+However, because these quantities are not antisymmetric under time reversal, it is unclear whether they appropriately quantify predictive causality.
+Instead, here we employ information flow, which is a pivotal concept within the framework of information thermodynamics [34, 35, 36, 37, 36, 38, 39] and has recently been applied to various systems, including standard fluid turbulence [40, 41, 42].
+Because the information flow is antisymmetric under time reversal and becomes zero in an equilibrium state where all probability currents are absent, it may reflect some aspects of the predictive causality consistent with the behavior of the probability currents.
 The constructed model can be interpreted as a standard two-variable predator-prey model that takes into account intrinsic fluctuations.
 We show that the model exhibits persistent fluctuating cyclic oscillations, called quasi-cycles [22, 23], due to stochastic amplification.
-The quasi-cycles can be observed for a wide range of parameters, including parameter values consistent with experiments [14] and gyrokinetic simulations [42].
+Quasi-cycles are distinguished from both deterministic and stochastic limit cycles by their shorter correlation time scales [43, 44].
+Furthermore, they can readily emerge in any system possessing a stable spiral fixed point subject to noise.
 This result suggests the possibility that the previously observed periodic oscillations in a toroidal plasma are not limit cycles, but rather quasi-cycles.
@@
 This information propagation occurs as long as zonal flow and turbulence coexist, regardless of the presence or absence of quasi-cycles.
-Thus, our information-theoretic analysis may provide a theoretical foundation for the regulation of turbulence through the control of zonal flow.
+Our finding suggests that turbulence behavior may be predictable to a certain extent based on zonal flow characteristics.
 Figure 1: (a) Schematic of the model.
@@
 The characteristic period is given by $2\pi/\omega_{c}\simeq 10$.
-Setup.
+II Setup
 […]
@@
 $W_{\rho}({\bm{n}}+{\bm{S}}_{\rho}|{\bm{n}})$ $=\Omega\rho\prod_{\alpha\in\{Z,W\}}\dfrac{1}{\Omega^{\nu^{\alpha}_{\rho}}}\dfrac{n^{\alpha}!}{(n^{\alpha}-\nu^{\alpha}_{\rho})!},$ (2)
-[…]
-System size expansion.
-—Since we are interested in the collective behavior of the system, we consider coarse-grained descriptions by taking the infinite volume limit $\Omega\rightarrow\infty$.
+where the stoichiometric coefficients ${\bm{\nu}}_{\rho}=(\nu^{Z}_{\rho},\nu^{W}_{\rho})$ denote the number of reactants Z and W involved in each reaction $\rho$, and $\Omega$ denotes the system size, which may be interpreted as a total volume of the plasma.
+Although we can estimate parameter values that are consistent with experiments and simulations, we consider this model as a qualitative description and will not explore its quantitative aspects in depth.
+We note that similar models have been used for the laminar-turbulence transition in standard fluids to investigate the connection with the directed percolation universality class [5, 6, 7, 8].
+III System size expansion
+Since we are interested in the collective behavior of the system, we consider coarse-grained descriptions by taking the infinite volume limit $\Omega\rightarrow\infty$.
 Here, we note that the two limits $t\rightarrow\infty$ and $\Omega\rightarrow\infty$ do not commute in this model, which is known as Keizer’s paradox [46, 47].
@@
 Below, we focus on the opposite case, where we first take $\Omega\rightarrow\infty$ and then $t\rightarrow\infty$.
-The coarse-grained dynamics in the limit $\Omega\rightarrow\infty$ can be systematically obtained by applying van Kampen’s system size expansion [46].
+The coarse-grained dynamics in the limit $\Omega\rightarrow\infty$ can be systematically obtained by applying the van Kampen system size expansion (also known as the linear noise approximation) [48].
+From the central limit theorem, we first note that $p_{t}({\bm{n}}/\Omega)$ can be reasonably approximated as a Gaussian with a finite width of order $\Omega^{-1/2}$.
 To separate the fluctuating part from the average part, we introduce a new stochastic variable ${\bm{r}}_{t}=(z_{t},w_{t})\in\mathbb{R}^{2}$ as ${\bm{n}}_{t}/\Omega={\bm{c}}_{t}+{\bm{r}}_{t}/\Omega^{1/2}$, where ${\bm{c}}_{t}=(c^{Z}_{t},c^{W}_{t})\in\mathbb{R}^{2}_{\geq 0}$ denotes the mean energy density to be determined.
 By substituting this expression into Eq.
-(1) and expanding in terms of $\Omega^{-1/2}$, we find that the leading order $O(\Omega^{1/2})$ yields the time evolution equation for the average part ${\bm{c}}_{t}$, which can be regarded as a coarse-grained dynamics of Eq.
-(1):
+(1) and expanding in terms of $\Omega^{-1/2}$, we can obtain the time evolution equations for ${\bm{c}}_{t}$ and ${\bm{r}}_{t}$, respectively.
+III.1 Deterministic rate equation
+The leading order $O(\Omega^{1/2})$ yields the time evolution equation for the average part ${\bm{c}}_{t}$, which can be regarded as a coarse-grained dynamics of Eq. (1):
 $\dfrac{d}{dt}c^{Z}_{t}$ $=pc^{Z}_{t}c^{W}_{t}-dc^{Z}_{t},$ (3)
 $\dfrac{d}{dt}c^{W}_{t}$ $=bc^{W}_{t}-\gamma(c^{W}_{t})^{2}-pc^{Z}_{t}c^{W}_{t}.$ (4)
-See Ref.
-[47] for the detailed derivation.
+See Appendix A for a detailed derivation.
 This equation corresponds to the standard two-variable predator-prey model proposed by Diamond et al.
@@
 While cyclic oscillations emerge when $\gamma=0$, these cycles are not limit cycles but structurally unstable.
-Note that the coarse-grained description [Eqs.
+III.2 Linear Langevin equation
+The coarse-grained description [Eqs.
 (3) and (4)] ignores intrinsic noise that is inevitably present in the original individual-level description (1).
@@
 To investigate the effect of the intrinsic noise on the coexistence state, we focus on the regime $bp>d\gamma$ and proceed to the subleading order $O(\Omega^{0})$ of the system size expansion by substituting $(c^{Z}_{\mathrm{ss}},c^{W}_{\mathrm{ss}})=((bp-d\gamma)/p^{2},d/p)$.
-The resulting equation describes the time evolution of the fluctuating part ${\bm{r}}_{t}$ around ${\bm{c}}_{t}={\bm{c}}_{\mathrm{ss}}$ [47]:
+The resulting equation describes the time evolution of the fluctuating part ${\bm{r}}_{t}$ around ${\bm{c}}_{t}={\bm{c}}_{\mathrm{ss}}$ (see Appendix A for a detailed derivation):
 $\dfrac{d}{d\tau}z_{\tau}$ $=\epsilon w_{\tau}+\zeta^{Z}_{\tau},$ (5)
@@
 The terms $\zeta^{Z}_{\tau}$ and $\zeta^{W}_{\tau}$ denote the zero-mean white Gaussian noise that satisfies $\langle\zeta^{Z}_{\tau}\zeta^{Z}_{\tau^{\prime}}\rangle=2D^{Z}\delta(\tau-\tau^{\prime})$, $\langle\zeta^{Z}_{\tau}\zeta^{W}_{\tau^{\prime}}\rangle=-D^{Z}\delta(\tau-\tau^{\prime})$, and $\langle\zeta^{W}_{\tau}\zeta^{W}_{\tau^{\prime}}\rangle=2D^{W}\delta(\tau-\tau^{\prime})$, where $D^{Z}:=\epsilon c^{W}_{\mathrm{ss}}=\epsilon d/p$ and $D^{W}:=c^{W}_{\mathrm{ss}}=d/p$.
-Note that this linear Langevin equation has a form in which noise is added to the linearized Lotka–Volterra equation around the coexistence fixed point.
-We emphasize that the noise intensities $D^{Z}$ and $D^{W}$ are uniquely determined from the master equation (1).
+We emphasize that the statistical properties of the noise, such as the noise intensities $D^{Z}$ and $D^{W}$, are uniquely determined from the master equation (1).
 In other words, the noise driving the system is not artificially added but rather arises from intrinsic stochasticity.
-This property marks a stark difference from the stochastic models used in Refs.
-[26, 27, 28, 29, 30, 31, 32], where noise is artificially added to deterministic predator-prey models.
-The Langevin equations (5) and (6) are equivalent to the Fokker–Planck equation $\partial_{\tau}p_{\tau}({\bm{r}})=-\partial_{z}J^{Z}_{\tau}({\bm{r}})-\partial_{w}J^{W}_{\tau}({\bm{r}})$, where $p_{\tau}({\bm{r}})$ denotes the probability density of finding ${\bm{r}}_{\tau}={\bm{r}}$ and $J^{Z}_{\tau}({\bm{r}})$ and $J^{W}_{\tau}({\bm{r}})$ denote the probability currents [47].
-Quasi-cycles induced by stochastic amplification.
-—Although the coarse-grained deterministic description [Eqs.(3) and (4)] cannot predict stable cycles, the stochastic description [Eqs.
+This property contrasts with the stochastic models proposed in Refs.
+[26, 27, 29, 30, 31, 32, 33], where noise is artificially added to the corresponding deterministic predator-prey models, and its properties can be arbitrarily determined.
+We also remark that the noises $\zeta^{Z}_{\tau}$ and $\zeta^{W}_{\tau}$ in our model are not independent but are correlated as $\langle\zeta^{Z}_{\tau}\zeta^{W}_{\tau^{\prime}}\rangle=-D^{Z}\delta(\tau-\tau^{\prime})$, which is a property not considered in previously proposed models.
+Note that the linear Langevin equations (5) and (6) have a form in which noise is added to the linearized Lotka–Volterra equation around the coexistence fixed point.
+To see this point explicitly, we rewrite Eqs.
+(5) and (6) in vector form with ${\bm{r}}_{\tau}=(z_{\tau},w_{\tau})$:
+$\dfrac{d}{d\tau}{\bm{r}}_{\tau}=-\mathsf{A}{\bm{r}}_{\tau}+\mathsf{B}{\bm{\xi}}_{\tau},$ (7)
+where
+$\mathsf{A}=\begin{pmatrix}0&-\epsilon\\ \beta&1-\epsilon\end{pmatrix}$, (8)
+and $\bm{\xi}$} denotes the zero-mean white Gaussian noise that satisfies $\langle\xi^{\alpha}_{\tau}\xi^{\alpha^{\prime}}_{\tau^{\prime}}\rangle=\delta^{\alpha\alpha^{\prime}}\delta(\tau-\tau^{\prime})$ ($\alpha=Z,W$), which is multiplied by the noise matrix $\mathsf{B}$} that satisfies
+$\mathsf{B}\mathsf{B}^{\top}=\begin{pmatrix}2D^{Z}&-D^{Z}\\ -D^{Z}&2D^{W}\end{pmatrix}$. (9)
+Then, it is easy to see that $\mathsf{A}$ corresponds to the coefficient matrix in the linearized Lotka–Volterra equation (3) and (4) at the coexistence fixed point.
+The linear Langevin equations (5) and (6) are equivalent to the Fokker–Planck equation,
+$\partial_{\tau}p_{\tau}({\bm{r}})=-\dfrac{\partial}{\partial z}J^{Z}_{\tau}({\bm{r}})-\dfrac{\partial}{\partial w}J^{W}_{\tau}({\bm{r}}),$ (10)
+where $p_{\tau}({\bm{r}})$ denotes the probability density of finding ${\bm{r}}_{\tau}={\bm{r}}$ and $J^{Z}_{\tau}({\bm{r}})$ and $J^{W}_{\tau}({\bm{r}})$ denote the probability currents defined by
+$J^{Z}_{\tau}({\bm{r}})$ $:=\epsilon wp_{\tau}({\bm{r}})-D^{Z}\dfrac{\partial}{\partial z}p_{\tau}({\bm{r}})+\dfrac{D^{Z}}{2}\dfrac{\partial}{\partial w}p_{\tau}({\bm{r}}),$ (11)
+$J^{W}_{\tau}({\bm{r}})$ $:=\bigl[-\beta z-(1-\epsilon)w\bigr]p_{\tau}({\bm{r}})$
+$\qquad-D^{W}\dfrac{\partial}{\partial w}p_{\tau}({\bm{r}})+\dfrac{D^{Z}}{2}\dfrac{\partial}{\partial z}p_{\tau}({\bm{r}}).$ (12)
+Because the coefficient matrix $\mathsf{A}$ has eigenvalues $\lambda_{\pm}$ with strictly positive real parts, there exists a unique steady-state probability density $p_{\mathrm{ss}}({\bm{r}})$ for this Fokker–Planck equation [49, 50].
+It can be explicitly calculated as
+$p_{\mathrm{ss}}({\bm{r}})=\dfrac{1}{2\pi\sqrt{\det\Sigma}}\exp\left(-\dfrac{1}{2}({\bm{r}}-\langle{\bm{r}}\rangle)^{\top}\Sigma^{-1}({\bm{r}}-\langle{\bm{r}}\rangle)\right),$ (13)
+where $\Sigma:=\langle({\bm{r}}-\langle{\bm{r}}\rangle)({\bm{r}}-\langle{\bm{r}}\rangle)^{\top}\rangle$ denotes the covariance matrix and $\langle\cdot\rangle$ denotes the average with respect to this distribution.
+Importantly, the steady state is generally out of equilibrium.
+That is, the steady-state probability current does not vanish, and there exists $\bm{r}$} such that ${\bm{J}}_{\mathrm{ss}}({\bm{r}}):=(J^{Z}_{\mathrm{ss}}({\bm{r}}),J^{W}_{\mathrm{ss}}({\bm{r}}))\neq{\bm{0}}$.
+In other words, the detailed balance condition (potential condition [49]) is generally violated in this linear Langevin equation.
+We can check that the detailed balance condition is satisfied when $\epsilon=0$.
+IV Quasi-cycles induced by stochastic amplification
+Although the coarse-grained deterministic description [Eqs.(3) and (4)] cannot predict stable cycles, the stochastic description [Eqs.
 (5) and (6)] predicts persistent cyclic oscillations called quasi-cycles.
@@
 Notably, the fluctuations of zonal flow follow those of turbulence with a phase lag $\sim\pi/2$, which are similar to the predator-prey-like oscillations observed in a toroidal plasma [14, 16, 12, 51, 52].
-This oscillatory behavior occurs when the coexistence state is a stable spiral fixed point in the coarse-grained deterministic description.
-More precisely, from the calculation of the power spectral density [47], we can show that the resonant stochastic amplification occurs for a wide range of parameters, where the characteristic resonant frequency $\omega_{c}:=\sqrt{\epsilon\beta-(1-\epsilon)^{2}/2}$ takes a real value (see also Fig.
+Here, it should be noted that several experimental studies have also observed the opposite behavior, where turbulence follows zonal flow [15, 17, 53].
+Whether such behavior can be reproduced in our model requires further investigation in future studies.
+This resonant behavior can be quantified by the power spectral density matrix, which is defined as the Fourier transform of the steady-state autocorrelation function [49]:
+$\mathsf{S}(\omega)$ $:=\dfrac{1}{2\pi}\int^{\infty}_{-\infty}\langle({\bm{r}}_{\tau}-\langle{\bm{r}}_{\tau}\rangle)({\bm{r}}_{0}-\langle{\bm{r}}_{0}\rangle)^{\top}\rangle e^{i\omega\tau}d\tau$
+$=\dfrac{1}{2\pi}\Bigl(\mathsf{A}-i\omega\mathsf{I}\Bigr)^{-1}\mathsf{B}\mathsf{B}^{\top}\Bigl(\mathsf{A}^{\top}+i\omega\mathsf{I}\Bigr)^{-1}.$ (14)
+Then, the power spectral density for zonal flow $Z$ can be calculated as
+$\mathsf{S}^{ZZ}(\omega)$ $=\dfrac{C_{1}+C_{2}\omega^{2}}{(\omega^{2}-\epsilon\beta)^{2}+(1-\epsilon)^{2}\omega^{2}},$ (15)
+where
+$C_{1}$ $:=\dfrac{1}{\pi}\Bigl((1-3\epsilon)D^{Z}+\epsilon^{2}D^{W}\Bigr),$ (16)
+$C_{2}$ $:=\dfrac{D^{Z}}{\pi}.$ (17)
+From this expression, we can see that the quasi-cycle occurs for a wide range of parameters where the characteristic resonant frequency $\omega_{c}:=\sqrt{\epsilon\beta-(1-\epsilon)^{2}/2}$ takes a real value (see also Fig.
 2(b)).
-Because the white noise in the Langevin equation covers all frequencies, the corresponding cycle with the resonant frequency $\omega_{c}$ is amplified.
+When this resonance condition is satisfied, the coexistence fixed point is a stable spiral fixed point with the characteristic frequency $|\mathrm{Im}[\lambda_{\pm}]|=\sqrt{\epsilon\beta-(1-\epsilon)^{2}/4}$, where $\mathrm{Im}[\cdot]$ denotes the imaginary part.
+Note that no tuning is necessary to achieve resonance in this model.
+Indeed, because the system is driven by the white noise, which covers all frequencies, the resonant frequency $\omega_{c}\sim|\mathrm{Im}[\lambda_{\pm}]|$ is excited without tuning.
 This mechanism is called stochastic amplification to distinguish it from stochastic resonance [22].
-The quasi-cycles differ from limit cycles in that they are stochastic oscillations where the different frequencies near $\omega_{c}$ are all excited [23, 48].
-Furthermore, because quasi-cycles can occur as long as the system has a stable spiral fixed point and noise, they may be widely observed in plasma turbulence under various conditions.
-Although we have emphasized that our model is a qualitative description, here we present a set of parameter values consistent with experiments and simulations for reference; By comparing with the gyrokinetic simulations [42], they can be estimated as $\epsilon\sim 1$ and $\beta\sim 10^{-4}$, which satisfies the resonance condition, and the corresponding cycle frequency reads $\omega_{c}\sim 10^{-2}$ in dimensionless units.
-Information propagation between zonal flow and turbulence.
-—The presence of quasi-cycles suggests that intrinsic fluctuations are essential for understanding the dynamics of zonal flow and turbulence.
-Therefore, we scrutinize the statistical causality underlying the intrinsic fluctuations from an information-theoretic viewpoint.
+We remark that a similar analysis can be carried out for turbulence $W$.
+Note that stochastic oscillations can also occur for a dynamical system with limit cycles perturbed by extrinsic noise.
+While quasi-cycles resemble such stochastic limit cycles, they possess several distinguishing characteristics.
+First, the power spectral density of quasi-cycles exhibits a long tail with $\omega^{-2}$ scaling [Eq.
+15]; in contrast, stochastic limit cycles decay more rapidly as $\omega^{-4}$ [44].
+This implies that the correlation time scale of fluctuations in quasi-cycles is shorter than that of stochastic limit cycles [43].
+Second, the joint probability density of the oscillating variables for quasi-cycles is Gaussian [Eq.
+(13)], whereas that for stochastic limit cycles takes the shape of a crater ridge [43].
+Furthermore, because quasi-cycles can readily emerge in any system possessing a stable spiral fixed point subject to noise, they may be more widely observed in plasma turbulence under various conditions.
+Although we have emphasized that our model is a qualitative description, here we present a set of parameter values consistent with experiments and simulations for reference.
+By comparing with the gyrokinetic simulations [54], they can be estimated as $\epsilon\sim 1$ and $\beta\sim 10^{-4}$, which satisfies the resonance condition, and the corresponding cycle frequency reads $\omega_{c}\sim 10^{-2}$ in dimensionless units.
+V Information propagation between zonal flow and turbulence
+The presence of quasi-cycles suggests that intrinsic fluctuations are essential for understanding the dynamics of zonal flow and turbulence.
+Therefore, we scrutinize the predictive causality underlying the intrinsic fluctuations from an information-theoretic viewpoint.
+In the context of plasma turbulence, (net) transfer entropy [24, 25] and information length [26, 27, 28, 29, 30, 31, 32, 33] have been used to quantify predictive causality.
+Although these quantities are useful in quantifying various aspects of causal influences, it should be noted that they are not antisymmetric under time reversal.
+In particular, the (net) transfer entropy can be nonzero even in an equilibrium state where all probability currents vanish [37].
+(For other shortcomings of the transfer entropy, see, e.g., Refs.
+[55, 56].) For the information length, although it becomes zero in such an equilibrium state, it also vanishes in a nonequilibrium steady state where some probability currents remain nonzero.
+Therefore, from the perspective of consistency with the behavior of the probability currents, it is unclear whether these quantities appropriately quantify predictive causality.
+In this study, we employ information flow [34, 35, 36, 37] because it is antisymmetric under time reversal and becomes zero in the equilibrium state.
 We first introduce the mutual information [57], which quantifies the mutual dependence between zonal flow and turbulence:
@@
 Figure 2: (a) Schematic of information flow $\dot{I}^{Z\rightarrow W}_{\lambda}$.
-(b) $(\epsilon,\beta)$-dependence of $\dot{I}^{Z\rightarrow W}_{\lambda=1/2}$.
+(b) $(\epsilon,\beta)$-dependence of $\dot{I}^{Z\rightarrow W}_{\lambda=1/2}$ with $D^{W}=1/2$.
 The dashed line denotes the resonance condition $2\epsilon\beta-(1-\epsilon)^{2}=0$.
@@
 Here, we note that the non-bipartite structure allows the case where the sign of $\dot{I}^{Z\rightarrow W}_{\lambda}$ varies depending on $\lambda$, which makes it difficult to determine the direction of information transfer [37].
-In that case, $\dot{I}^{Z\rightarrow W}_{\lambda=1/2}$ plays a special role because it is odd under time reversal and vanishes in the equilibrium state.
+In that case, $\dot{I}^{Z\rightarrow W}_{\lambda=1/2}$ plays a special role because it is antisymmetric under time reversal and vanishes in the equilibrium state.
 Fortunately, as will be shown below, this problem is irrelevant to our model.
@@
 where $p_{\tau}(z)$ and $p_{\tau}(w)$ are marginal distributions for zonal flow and turbulence, respectively.
-By noting that the steady-state distribution $p_{\mathrm{ss}}({\bm{r}})$ for the linear Langevin equations (5) and (6) is Gaussian with covariance matrix $\Sigma:=\langle({\bm{r}}-\langle{\bm{r}}\rangle)({\bm{r}}-\langle{\bm{r}}\rangle)^{\top}\rangle$, the information flow in the steady state can be calculated as
+By noting that the steady-state distribution $p_{\mathrm{ss}}({\bm{r}})$ is Gaussian [Eq.
+(13)], the information flow in the steady state can be calculated as
 $\dot{I}^{Z\rightarrow W}_{\lambda}$ $=\dfrac{D^{Z}}{\det\Sigma}\left[\Sigma^{WW}+(1-\lambda)\Sigma^{ZW}\right]$
@@
 where $\Sigma^{ZW}=\langle(z-\langle z\rangle)(w-\langle w\rangle)\rangle$ and $\Sigma^{WW}=\langle(w-\langle w\rangle)^{2}\rangle$.
-See Ref.
-[47] for the detailed derivation of Eqs.
+See Appendix B for the detailed derivation of Eqs.
 (20) and (21).
@@
 2(b) suggests that the behavior of the information flow does not depend significantly on whether stochastic amplification occurs or not.
-In other words, even when the system does not exhibit the fluctuating predator-prey oscillation, there is a statistical causality from zonal flow to turbulence.
+In other words, even when the system does not exhibit the fluctuating predator-prey oscillation, there is a predictive causality from zonal flow to turbulence.
 This information-theoretic analysis may provide a theoretical foundation for the regulation of turbulence through the control of zonal flow.
-We finally remark that the inferred causality from the information flow is consistent with that from the cross-correlation function, although interpreting the latter is less straightforward [47].
-Concluding remarks.
-—We have proposed a simple stochastic predator-prey model for plasma turbulence that exhibits quasi-cycles induced by stochastic amplification of intrinsic noise.
-For this model, we have quantified the causality between zonal flow and turbulence by showing that the information of zonal flow is propagated to turbulence.
-Our model illustrates the significance of considering the effects of intrinsic noise in understanding the predator-prey-like oscillations observed in plasma turbulence.
+We finally remark that the inferred causality from the information flow is consistent with that from the cross-correlation function, although interpreting the latter is less straightforward (see Appendix C).
+VI Concluding remarks
+We have proposed a simple stochastic predator-prey model for plasma turbulence that exhibits quasi-cycles induced by stochastic amplification of intrinsic noise.
+For this model, we have quantified the predictive causality between zonal flow and turbulence by showing that the information of zonal flow is propagated to turbulence.
+Several remarks on our results are in order.
+First, although we have considered the four simple “reactions” that incorporate a minimal self-regulation mechanism, it is straightforward to extend our approach to more complicated reactions, such as those involving time-dependent rate constants or additional reactant species (e.g., the mean shear flow).
+Even in such a case, both the system size expansion and the information-theoretic analysis are applicable.
+Second, our model illustrates the significance of considering the effects of intrinsic noise in understanding the predator-prey-like oscillations observed in plasma turbulence.
 Notably, our results suggest the possibility that the observed cyclic oscillations in a toroidal plasma are quasi-cycles rather than limit cycles.
 Using the methods proposed in Ref.
-[48], it would be possible to distinguish quasi-cycles and limit cycles both in experiments and simulations.
+[43], it would be possible to distinguish quasi-cycles and (stochastic) limit cycles both in experiments and simulations.
 Further studies are needed to verify whether both quasi-cycles and information transfer from zonal flow to turbulence can be observed in experiments and simulations of more realistic models, such as the Hasegawa–Wakatani equation [58].
+Third, our results essentially hold true for general predator-prey dynamics, not limited to plasma turbulence.
+In particular, while the essence of our first main result—the emergence of quasi-cycles due to stochastic amplification in the predator-prey dynamics—has already been recognized in the field of ecology [22, 23], our second main result concerning information flow may offer new perspectives even within an ecological context.
+Finally, while we have shown that information of zonal flow is propagated to turbulence, further research is needed to understand how this result relates to the regulation of turbulence through the control of zonal flow.
+This issue also seems connected to the general question of how predictive causality relates to dynamical causality, which can be quantified in frameworks such as convergent cross mapping [59, 60].
+We hope our research will serve as a starting point for investigating these issues.
 Acknowledgements.
@@
 was supported by JSPS KAKENHI Grant Number JP25K00986, Japan.
+Appendix A Detailed calculation of system size expansion
+[…]
+[…]
+$=\rho(\tilde{c}^{Z})^{\nu^{Z}_{\rho}}(\tilde{c}^{W})^{\nu^{W}_{\rho}}.$ (22)
+Then, by taking the Kramers–Moyal expansion [49], the master equation [Eq. (1)] can be rewritten as
+[…]
+[…]
+$p_{t}(\tilde{\bm{c}})\asymp e^{-\Omega I_{t}(\tilde{\bm{c}})}$, (24)
+[…]
+$\tilde{\bm{c}}={\bm{c}}+\dfrac{\bm{r}}{\Omega^{1/2}},$ (25)
+where the mean-field density ${\bm{c}}\in\mathbb{R}^{2}_{\geq 0}$ is a function of time $t$ to be determined.
+We denote by ${\bm{c}}_{t}$ and ${\bm{r}}_{t}$ the value of $\bm{c}$} and $\bm{r}$} at time $t$, respectively.
+Correspondingly, let $p_{t}({\bm{r}}):=\Omega^{-1/2}p_{t}({\bm{c}}_{t}+\Omega^{-1/2}{\bm{r}})$ be the probability density of finding ${\bm{r}}_{t}={\bm{r}}$.
+Then, Eq.
+(23) can be expressed as
+$\partial_{t}p_{t}({\bm{r}})-\Omega^{1/2}\dot{\bm{c}}_{t}\cdot\dfrac{\partial}{\partial{\bm{r}}}p_{t}({\bm{r}})$
+$=\sum_{\rho}\sum^{\infty}_{k=1}\dfrac{\Omega^{1-k/2}}{k!}\left(-{\bm{S}}_{\rho}\cdot\dfrac{\partial}{\partial{\bm{r}}}\right)^{k}\mathcal{W}_{\rho}\left({\bm{c}}_{t}+\dfrac{\bm{r}}{\Omega^{1/2}}\right)p_{t}({\bm{r}}).$ (26)
+A.1 Deterministic rate equation
+The leading order $O(\Omega^{1/2})$ of Eq. (26) yields the deterministic rate equation for the average part ${\bm{c}}_{t}=(c^{Z}_{t},c^{W}_{t})$:
+$\dfrac{d}{dt}{\bm{c}}_{t}=\sum_{\rho}{\bm{S}}_{\rho}\mathcal{W}_{\rho}({\bm{c}}_{t}).$ (27)
+By noting that $\mathcal{W}_{d}({\bm{c}})=dc^{Z},\mathcal{W}_{b}({\bm{c}})=bc^{W},\mathcal{W}_{p}({\bm{c}})=pc^{Z}c^{W}$, and $\mathcal{W}_{\gamma}({\bm{c}})=\gamma(c^{W})^{2}$, the rate equation can be expressed as
+$\dfrac{d}{dt}c^{Z}_{t}$ $=pc^{Z}_{t}c^{W}_{t}-dc^{Z}_{t},$ (28)
+$\dfrac{d}{dt}c^{W}_{t}$ $=bc^{W}_{t}-\gamma(c^{W}_{t})^{2}-pc^{Z}_{t}c^{W}_{t},$ (29)
+which corresponds to Eqs. (3) and (4).
+A.2 Linear Langevin equation
+The subleading order $O(1)$ of Eq. (26) yields the Fokker–Planck equation:
+$\partial_{t}p_{t}({\bm{r}})=-\sum_{\rho}\left.\dfrac{\partial}{\partial{\bm{c}}}\mathcal{W}_{\rho}({\bm{c}})\right|_{{\bm{c}}={\bm{c}}_{t}}\cdot\left({\bm{S}}_{\rho}\cdot\dfrac{\partial}{\partial{\bm{r}}}\right){\bm{r}}p_{t}({\bm{r}})$
+$\qquad\qquad\quad+\sum_{\rho}\dfrac{1}{2}\mathcal{W}_{\rho}({\bm{c}}_{t})\left({\bm{S}}_{\rho}\cdot\dfrac{\partial}{\partial{\bm{r}}}\right)^{2}p_{t}({\bm{r}}).$ (30)
+Note that ${\bm{c}}_{t}$ in Eq.
+(30) obeys the deterministic rate equations (28) and (29).
+In other words, the nonlinearity of the predator-prey dynamics is reflected in the Fokker–Planck equation through the dependence on ${\bm{c}}_{t}$.
+If we substitute the fixed point solution ${\bm{c}}_{t}={\bm{c}}_{\mathrm{ss}}$, then the Fokker–Planck equation (30) describes the time evolution of the fluctuations around the fixed point.
+Since we are interested in the state where zonal flow and turbulence coexist, we focus on the regime $bp>d\gamma$ and substitute the stable coexistence fixed point $(c^{Z}_{\mathrm{ss}},c^{W}_{\mathrm{ss}})=((bp-d\gamma)/p^{2},d/p)$ into Eq.
+(30).
+By introducing the dimensionless time $\tau:=bt$ and dimensionless parameters $\beta:=d/b>0$ and $\epsilon:=1-d\gamma/bp$ ($0<\epsilon<1$), the resulting Fokker–Planck equation can be expressed as
+$\partial_{\tau}p_{\tau}({\bm{r}})=-\dfrac{\partial}{\partial z}J^{Z}_{\tau}({\bm{r}})-\dfrac{\partial}{\partial w}J^{W}_{\tau}({\bm{r}}),$ (31)
+where $J^{Z}_{\tau}({\bm{r}})$ and $J^{W}_{\tau}({\bm{r}})$ denote the dimensionless probability currents defined by
+$J^{Z}_{\tau}({\bm{r}})$ $:=\epsilon wp_{\tau}({\bm{r}})-D^{Z}\dfrac{\partial}{\partial z}p_{\tau}({\bm{r}})+\dfrac{D^{Z}}{2}\dfrac{\partial}{\partial w}p_{\tau}({\bm{r}}),$ (32)
+$J^{W}_{\tau}({\bm{r}})$ $:=\bigl[-\beta z-(1-\epsilon)w\bigr]p_{\tau}({\bm{r}})$
+$\qquad-D^{W}\dfrac{\partial}{\partial w}p_{\tau}({\bm{r}})+\dfrac{D^{Z}}{2}\dfrac{\partial}{\partial z}p_{\tau}({\bm{r}}),$ (33)
+with $D^{Z}:=\epsilon c^{W}_{\mathrm{ss}}=\epsilon d/p$ and $D^{W}:=c^{W}_{\mathrm{ss}}=d/p$. The Langevin equation corresponding to this Fokker–Planck equation reads
+$\dfrac{d}{d\tau}z_{\tau}$ $=\epsilon w_{\tau}+\zeta^{Z}_{\tau},$ (34)
+$\dfrac{d}{d\tau}w_{\tau}$ $=-\beta z_{\tau}-(1-\epsilon)w_{\tau}+\zeta^{W}_{\tau},$ (35)
+where $\zeta^{Z}_{\tau}$ and $\zeta^{W}_{\tau}$ denote the zero-mean white Gaussian noise that satisfies
+$\langle\zeta^{Z}_{\tau}\zeta^{Z}_{\tau^{\prime}}\rangle$ $=2D^{Z}\delta(\tau-\tau^{\prime}),$ (36)
+$\langle\zeta^{Z}_{\tau}\zeta^{W}_{\tau^{\prime}}\rangle$ $=-D^{Z}\delta(\tau-\tau^{\prime}),$ (37)
+$\langle\zeta^{W}_{\tau}\zeta^{W}_{\tau^{\prime}}\rangle$ $=2D^{W}\delta(\tau-\tau^{\prime}).$ (38)
+Note that $\beta$ and $\epsilon$ can be interpreted as a dimensionless damping rate and the ratio of time scales for $z$ and $w$, respectively.
+Here, we emphasize that the noise intensities $D^{Z}$ and $D^{W}$ are determined from the rate constants $\rho$ in the master equation [Eq.
+(1)].
+In other words, the noise that drives the system is not external but arises from the intrinsic stochasticity.
+We finally remark that the validity of the system size expansion up to the subleading order has been well confirmed in the literature.
+For example, in Ref.
+[61], it is reported that the linear Langevin equation agrees well with the master equation for sufficiently large system size.
+Appendix B Detailed calculation of information flow
+B.1 Derivation of Eq. (20)
+In this section, we derive Eq. (20). From the definition of $\dot{I}^{Z\rightarrow W}_{\lambda}$, we first note that
+$\dot{I}^{Z\rightarrow W}_{\lambda}$ $:=\lim_{\Delta\tau\rightarrow 0^{+}}\dfrac{I[Z_{\tau+\lambda\Delta\tau}\colon\!W_{\tau+\Delta\tau}]-I[Z_{\tau+\lambda\Delta\tau}\colon\!W_{\tau}]}{\Delta\tau}$
+$=\lim_{h\rightarrow 0^{+}}\dfrac{1}{h}\left(\int dzdwp(z,\tau+\lambda h;w,\tau+h)\ln\dfrac{p(z,\tau+\lambda h;w,\tau+h)}{p_{\tau+\lambda h}(z)p_{\tau+h}(w)}-\int dzdwp(z,\tau+\lambda h;w,\tau)\ln\dfrac{p(z,\tau+\lambda h;w,\tau)}{p_{\tau+\lambda h}(z)p_{\tau}(w)}\right),$ (39)
+where $p(z,\tau+\lambda h;w,\tau+h)$ and $p(z,\tau+\lambda h;w,\tau)$ denote the two-point probability densities.
+When $h=0$, the two-point probability densities correspond to the joint probability density at time $\tau$: $p(z,\tau;w,\tau)=p_{\tau}(z,w)$.
+By expanding $p(z,\tau+\lambda h;w,\tau+h)$, $p(z,\tau+\lambda h;w,\tau)$, $p_{\tau+h}(w)$, and $p_{\tau+\lambda h}(z)$ with respect to $h$, we obtain
+$\dot{I}^{Z\rightarrow W}_{\lambda}$ $=\int dz^{\prime}dw\left.\dfrac{d}{dh}p(z^{\prime},\tau+\lambda h;w,\tau+h)\right|_{h=0}\ln\dfrac{p_{\tau}(z^{\prime},w)}{p_{\tau}(z^{\prime})p_{\tau}(w)}$
+$\qquad-\int dzdw^{\prime}\left.\dfrac{d}{dh}p(z,\tau+\lambda h;w^{\prime},\tau)\right|_{h=0}\ln\dfrac{p_{\tau}(z,w^{\prime})}{p_{\tau}(z)p_{\tau}(w^{\prime})}.$ (40)
+Here, for later convenience, we have replaced $z$ and $w$ with $z^{\prime}$ and $w^{\prime}$ in the first and second integrals, respectively. By noting that
+$p(z^{\prime},\tau+\lambda h;w,\tau+h)-p(z^{\prime},\tau;w,\tau)$ $=p(z^{\prime},\tau+\lambda h;w,\tau+\lambda h+(1-\lambda)h)-p(z^{\prime},\tau+\lambda h;w,\tau+\lambda h)$
+$\qquad+p(z^{\prime},\tau+\lambda h;w,\tau+\lambda h)-p(z^{\prime},\tau;w,\tau),$ (41)
+the derivative $\frac{d}{dh}p(z^{\prime},\tau+\lambda h;w,\tau+h)|_{h=0}$ can be expressed as
+$\left.\dfrac{d}{dh}p(z^{\prime},\tau+\lambda h;w,\tau+h)\right|_{h=0}=(1-\lambda)\left.\dfrac{d}{dh}p(z^{\prime},\tau;w,\tau+h)\right|_{h=0}+\lambda\left.\dfrac{d}{dh}p_{\tau+h}(z^{\prime},w)\right|_{h=0}.$ (42)
+By substituting this expression into Eq. (40), we have
+$\dot{I}^{Z\rightarrow W}_{\lambda}$ $=(1-\lambda)\int dz^{\prime}dw\left.\dfrac{d}{dh}p(z^{\prime},\tau;w,\tau+h)\right|_{h=0}\ln\dfrac{p_{\tau}(z^{\prime},w)}{p_{\tau}(z^{\prime})p_{\tau}(w)}$
+$\qquad+\lambda\int dz^{\prime}dw\left.\dfrac{d}{dh}p_{\tau+h}(z^{\prime},w)\right|_{h=0}\ln\dfrac{p_{\tau}(z^{\prime},w)}{p_{\tau}(z^{\prime})p_{\tau}(w)}$
+$\qquad-\lambda\int dzdw^{\prime}\left.\dfrac{d}{dh}p(z,\tau+h;w^{\prime},\tau)\right|_{h=0}\ln\dfrac{p_{\tau}(z,w^{\prime})}{p_{\tau}(z)p_{\tau}(w^{\prime})}$
+$=(1-\lambda)\int dz^{\prime}dw^{\prime}dzdw\left.\dfrac{d}{dh}p(z,w,\tau+h|z^{\prime},w^{\prime},\tau)\right|_{h=0}p_{\tau}(z^{\prime},w^{\prime})\ln\dfrac{p_{\tau}(z^{\prime},w)}{p_{\tau}(z^{\prime})p_{\tau}(w)}$
+$\qquad+\lambda\int dzdw\left.\dfrac{d}{dh}p_{\tau+h}(z,w)\right|_{h=0}\ln\dfrac{p_{\tau}(z,w)}{p_{\tau}(z)p_{\tau}(w)}$
+$\qquad-\lambda\int dz^{\prime}dw^{\prime}dzdw\left.\dfrac{d}{dh}p(z,w,\tau+h|z^{\prime},w^{\prime},\tau)\right|_{h=0}p_{\tau}(z^{\prime},w^{\prime})\ln\dfrac{p_{\tau}(z,w^{\prime})}{p_{\tau}(z)p_{\tau}(w^{\prime})}.$ (43)
+In the second equality, we have used
+$p(z^{\prime},\tau;w,\tau+h)$ $=\int dw^{\prime}dzp(z^{\prime},w^{\prime},\tau;z,w,\tau+h)=\int dw^{\prime}dzp(z,w,\tau+h|z^{\prime},w^{\prime},\tau)p_{\tau}(z^{\prime},w^{\prime}),$ (44)
+$p(z,\tau+h;w^{\prime},\tau)$ $=\int dwdz^{\prime}p(z,w,\tau+h;z^{\prime},w^{\prime},\tau)=\int dwdz^{\prime}p(z,w,\tau+h|z^{\prime},w^{\prime},\tau)p_{\tau}(z^{\prime},w^{\prime}),$ (45)
+where $p(z,w,\tau+h|z^{\prime},w^{\prime},\tau)$ denotes the conditional probability density. Note that this conditional probability density obeys the Fokker–Planck equation (31):
+$\left.\dfrac{d}{dh}p(z,w,\tau+h|z^{\prime},w^{\prime},\tau)\right|_{h=0}$ $=-\partial_{z}\left[F^{Z}({\bm{r}})\delta({\bm{r}}-{\bm{r}}^{\prime})-D^{Z}\partial_{z}\delta({\bm{r}}-{\bm{r}}^{\prime})+\dfrac{D^{Z}}{2}\partial_{w}\delta({\bm{r}}-{\bm{r}}^{\prime})\right]$
+$\qquad-\partial_{w}\left[F^{W}({\bm{r}})\delta({\bm{r}}-{\bm{r}}^{\prime})-D^{W}\partial_{w}\delta({\bm{r}}-{\bm{r}}^{\prime})+\dfrac{D^{Z}}{2}\partial_{z}\delta({\bm{r}}-{\bm{r}}^{\prime})\right],$ (46)
+where $F^{Z}({\bm{r}}):=\epsilon w$ and $F^{W}({\bm{r}}):=-\beta z-(1-\epsilon)w$, and we have used $p(z,w,\tau|z^{\prime},w^{\prime},\tau)=\delta({\bm{r}}-{\bm{r}}^{\prime})$.
+Then, Eq.
+(43) can be calculated as
+$\dot{I}^{Z\rightarrow W}_{\lambda}$ $=-(1-\lambda)\int dz^{\prime}dw^{\prime}dzdw\partial_{z}\left[F^{Z}({\bm{r}})\delta({\bm{r}}-{\bm{r}}^{\prime})-D^{Z}\partial_{z}\delta({\bm{r}}-{\bm{r}}^{\prime})+\dfrac{D^{Z}}{2}\partial_{w}\delta({\bm{r}}-{\bm{r}}^{\prime})\right]p_{\tau}(z^{\prime},w^{\prime})\ln\dfrac{p_{\tau}(z^{\prime},w)}{p_{\tau}(z^{\prime})p_{\tau}(w)}$
+$\quad-(1-\lambda)\int dz^{\prime}dw^{\prime}dzdw\partial_{w}\left[F^{W}({\bm{r}})\delta({\bm{r}}-{\bm{r}}^{\prime})-D^{W}\partial_{w}\delta({\bm{r}}-{\bm{r}}^{\prime})+\dfrac{D^{Z}}{2}\partial_{z}\delta({\bm{r}}-{\bm{r}}^{\prime})\right]p_{\tau}(z^{\prime},w^{\prime})\ln\dfrac{p_{\tau}(z^{\prime},w)}{p_{\tau}(z^{\prime})p_{\tau}(w)}$
+$\quad+\lambda\int dzdw\left[-\partial_{z}J^{Z}_{\tau}({\bm{r}})-\partial_{w}J^{W}_{\tau}({\bm{r}})\right]\ln\dfrac{p_{\tau}(z,w)}{p_{\tau}(z)p_{\tau}(w)}$
+$\quad+\lambda\int dz^{\prime}dw^{\prime}dzdw\partial_{z}\left[F^{Z}({\bm{r}})\delta({\bm{r}}-{\bm{r}}^{\prime})-D^{Z}\partial_{z}\delta({\bm{r}}-{\bm{r}}^{\prime})+\dfrac{D^{Z}}{2}\partial_{w}\delta({\bm{r}}-{\bm{r}}^{\prime})\right]p_{\tau}(z^{\prime},w^{\prime})\ln\dfrac{p_{\tau}(z,w^{\prime})}{p_{\tau}(z)p_{\tau}(w^{\prime})}$
+$\quad+\lambda\int dz^{\prime}dw^{\prime}dzdw\partial_{w}\left[F^{W}({\bm{r}})\delta({\bm{r}}-{\bm{r}}^{\prime})-D^{W}\partial_{w}\delta({\bm{r}}-{\bm{r}}^{\prime})+\dfrac{D^{Z}}{2}\partial_{z}\delta({\bm{r}}-{\bm{r}}^{\prime})\right]p_{\tau}(z^{\prime},w^{\prime})\ln\dfrac{p_{\tau}(z,w^{\prime})}{p_{\tau}(z)p_{\tau}(w^{\prime})}$
+$=-(1-\lambda)\int dzdw\partial_{z}\left[F^{Z}({\bm{r}})p_{\tau}(z,w)\ln\dfrac{p_{\tau}(z,w)}{p_{\tau}(z)p_{\tau}(w)}-D^{Z}\partial_{z}\left(p_{\tau}(z,w)\ln\dfrac{p_{\tau}(z,w)}{p_{\tau}(z)p_{\tau}(w)}\right)\right$.
+$\qquad\left.+\dfrac{D^{Z}}{2}\biggl(\partial_{w}p_{\tau}(z,w)\biggr)\ln\dfrac{p_{\tau}(z,w)}{p_{\tau}(z)p_{\tau}(w)}\right]$
+$\quad-(1-\lambda)\int dzdw\left\{\partial_{w}\left[F^{W}({\bm{r}})p_{\tau}(z,w)-D^{W}\partial_{w}p_{\tau}(z,w)\right]\ln\dfrac{p_{\tau}(z,w)}{p_{\tau}(z)p_{\tau}(w)}+\dfrac{D^{Z}}{2}\partial_{z}\left[\biggl(\partial_{w}p_{\tau}(z,w)\biggr)\ln\dfrac{p_{\tau}(z,w)}{p_{\tau}(z)p_{\tau}(w)}\right]\right\}$
+$\quad+\lambda\int dzdw\left[-\partial_{z}J^{Z}_{\tau}({\bm{r}})-\partial_{w}J^{W}_{\tau}({\bm{r}})\right]\ln\dfrac{p_{\tau}(z,w)}{p_{\tau}(z)p_{\tau}(w)}$
+$\quad+\lambda\int dzdw\left\{\partial_{z}\left[F^{Z}({\bm{r}})p_{\tau}(z,w)-D^{Z}\partial_{z}p_{\tau}(z,w)\right]\ln\dfrac{p_{\tau}(z,w)}{p_{\tau}(z)p_{\tau}(w)}+\dfrac{D^{Z}}{2}\partial_{w}\left[\biggl(\partial_{z}p_{\tau}(z,w)\biggr)\ln\dfrac{p_{\tau}(z,w)}{p_{\tau}(z)p_{\tau}(w)}\right]\right\}$
+$\quad+\lambda\int dzdw\partial_{w}\left[F^{W}({\bm{r}})p_{\tau}(z,w)\ln\dfrac{p_{\tau}(z,w)}{p_{\tau}(z)p_{\tau}(w)}-D^{W}\partial_{w}\left(p_{\tau}(z,w)\ln\dfrac{p_{\tau}(z,w)}{p_{\tau}(z)p_{\tau}(w)}\right)\right$.
+$\qquad\left.+\dfrac{D^{Z}}{2}\biggl(\partial_{z}p_{\tau}(z,w)\biggr)\ln\dfrac{p_{\tau}(z,w)}{p_{\tau}(z)p_{\tau}(w)}\right].$ (47)
+Note that, in the second equality, the first and last integrals and the last terms on both the second and fourth integrals vanish, assuming the natural boundary condition. Then, we have
+$\dot{I}^{Z\rightarrow W}_{\lambda}$ $=-(1-\lambda)\int dzdw\partial_{w}\left[F^{W}({\bm{r}})p_{\tau}(z,w)-D^{W}\partial_{w}p_{\tau}(z,w)\right]\ln\dfrac{p_{\tau}(z,w)}{p_{\tau}(z)p_{\tau}(w)}$
+$\quad+\lambda\int dzdw\left[-\partial_{z}J^{Z}_{\tau}({\bm{r}})-\partial_{w}J^{W}_{\tau}({\bm{r}})\right]\ln\dfrac{p_{\tau}(z,w)}{p_{\tau}(z)p_{\tau}(w)}$
+$\quad+\lambda\int dzdw\partial_{z}\left[F^{Z}({\bm{r}})p_{\tau}(z,w)-D^{Z}\partial_{z}p_{\tau}(z,w)\right]\ln\dfrac{p_{\tau}(z,w)}{p_{\tau}(z)p_{\tau}(w)}.$ (48)
+By noting that
+$J^{Z}_{\tau}({\bm{r}})$ $:=F^{Z}({\bm{r}})p_{\tau}({\bm{r}})-D^{Z}\partial_{z}p_{\tau}({\bm{r}})+\dfrac{D^{Z}}{2}\partial_{w}p_{\tau}({\bm{r}}),$ (49)
+$J^{W}_{\tau}({\bm{r}})$ $:=F^{W}({\bm{r}})p_{\tau}({\bm{r}})-D^{W}\partial_{w}p_{\tau}({\bm{r}})+\dfrac{D^{Z}}{2}\partial_{z}p_{\tau}({\bm{r}}),$ (50)
+we obtain
+$\dot{I}^{Z\rightarrow W}_{\lambda}$ $=-\int dzdw\partial_{w}\left[F^{W}({\bm{r}})p_{\tau}(z,w)-D^{W}\partial_{w}p_{\tau}(z,w)+\lambda D^{Z}\partial_{z}p_{\tau}(z,w)\right]\ln\dfrac{p_{\tau}(z,w)}{p_{\tau}(z)p_{\tau}(w)}$
+$=\int dzdw\left[J^{W}_{\tau}({\bm{r}})+(2\lambda-1)\dfrac{D^{Z}}{2}\partial_{z}p_{\tau}({\bm{r}})\right]\partial_{w}\ln\dfrac{p_{\tau}({\bm{r}})}{p_{\tau}(z)p_{\tau}(w)}.$ (51)
+We thus arrive at Eq. (20). Similarly, we can also derive the following expression for $\dot{I}^{W\rightarrow Z}_{1-\lambda}$:
+$\dot{I}^{W\rightarrow Z}_{1-\lambda}=\int dzdw\left[J^{Z}_{\tau}({\bm{r}})-(2\lambda-1)\dfrac{D^{Z}}{2}\partial_{w}p_{\tau}({\bm{r}})\right]\partial_{z}\ln\dfrac{p_{\tau}({\bm{r}})}{p_{\tau}(z)p_{\tau}(w)}.$ (52)
+We can easily check that Eqs.
+(51) and (52) satisfy $d_{\tau}I[Z\colon\!W]=\dot{I}^{Z\rightarrow W}_{\lambda}+\dot{I}^{W\rightarrow Z}_{1-\lambda}$.
+Finally, we remark that $\dot{I}^{Z\rightarrow W}_{\lambda}$ can also be calculated from $\dot{I}^{Z\rightarrow W}_{\lambda=0}$ and $\dot{I}^{Z\rightarrow W}_{\lambda=1}$ as follows [37]:
+$\dot{I}^{Z\rightarrow W}_{\lambda}=(1-\lambda)\dot{I}^{Z\rightarrow W}_{\lambda=0}+\lambda\dot{I}^{Z\rightarrow W}_{\lambda=1}.$ (53)
+B.2 Derivation of Eq. (21)
+In this section, we derive Eq.
+(21).
+The stationary joint distribution $p_{\mathrm{ss}}({\bm{r}})$ and marginal distributions $p_{\mathrm{ss}}(z)$ and $p_{\mathrm{ss}}(w)$ can be explicitly calculated as [49]
+$p_{\mathrm{ss}}({\bm{r}})$ $=\dfrac{1}{2\pi\sqrt{\det\Sigma}}\exp\left(-\dfrac{1}{2}({\bm{r}}-\langle{\bm{r}}\rangle)^{\top}\Sigma^{-1}({\bm{r}}-\langle{\bm{r}}\rangle)\right),$ (54)
+$p_{\mathrm{ss}}(z)$ $=\dfrac{1}{\sqrt{2\pi\Sigma^{ZZ}}}\exp\left(-\dfrac{1}{2}\dfrac{(z-\langle z\rangle)^{2}}{\Sigma^{ZZ}}\right),$ (55)
+$p_{\mathrm{ss}}(w)$ $=\dfrac{1}{\sqrt{2\pi\Sigma^{WW}}}\exp\left(-\dfrac{1}{2}\dfrac{(w-\langle w\rangle)^{2}}{\Sigma^{WW}}\right),$ (56)
+where $\Sigma$ denotes the covariance matrix
+$\Sigma=\begin{pmatrix}\Sigma^{ZZ}&\Sigma^{ZW}\\ \Sigma^{WZ}&\Sigma^{WW}\end{pmatrix}=\begin{pmatrix}\langle(z-\langle z\rangle)^{2}\rangle&\langle(z-\langle z\rangle)(w-\langle w\rangle)\rangle\\ \langle(z-\langle z\rangle)(w-\langle w\rangle)\rangle&\langle(w-\langle w\rangle)^{2}\rangle\end{pmatrix}$, (57)
+which satisfies the Lyapunov equation:
+$\mathsf{A}\Sigma+\Sigma\mathsf{A}^{\top}=\mathsf{B}\mathsf{B}^{\top},$ (58)
+where $\mathsf{A}$ and $\mathsf{B}$ are matrices defined by Eqs. (8) and (9). By solving this equation, the covariance matrix can be calculated as [49]
+$\Sigma$ $=\dfrac{(\det\mathsf{A})\mathsf{B}\mathsf{B}^{\top}+(\mathsf{A}-(\mathop{\mathrm{tr}}\mathsf{A})\mathsf{I})\mathsf{B}\mathsf{B}^{\top}(\mathsf{A}-(\mathop{\mathrm{tr}}\mathsf{A})\mathsf{I})^{\top}}{2(\mathop{\mathrm{tr}}\mathsf{A})(\det\mathsf{A})}$
+$=\dfrac{1}{2\beta\epsilon(1-\epsilon)}\begin{pmatrix}2D^{Z}\left[\beta\epsilon+(1-\epsilon)(2(1-\epsilon)-1)\right]+2D^{W}\epsilon^{2}&-2\beta(1-\epsilon)D^{Z}\\ -2\beta(1-\epsilon)D^{Z}&2D^{W}\beta\epsilon+2\beta^{2}D^{Z}\end{pmatrix}$.
+(59)
+To derive Eq.
+(21), it is easy to calculate $\dot{I}^{W\rightarrow Z}_{1-\lambda}$ instead of $\dot{I}^{Z\rightarrow W}_{\lambda}$ by using the relation $\dot{I}^{W\rightarrow Z}_{1-\lambda}=-\dot{I}^{Z\rightarrow W}_{\lambda}$ in the steady state.
+By substituting (49), (54), and (55) into (52), we can explicitly calculate the information flow as follows:
+$\dot{I}^{Z\rightarrow W}_{\lambda}=-\dot{I}^{W\rightarrow Z}_{1-\lambda}$ $=-\int dzdw\left(\epsilon wp_{\mathrm{ss}}({\bm{r}})-D^{Z}\partial_{z}p_{\mathrm{ss}}({\bm{r}})+(1-\lambda)D^{Z}\partial_{w}p_{\mathrm{ss}}({\bm{r}})\right)$
+$\qquad\times\left(-\dfrac{\Sigma^{WW}}{\det\Sigma}(z-\langle z\rangle)+\dfrac{\Sigma^{ZW}}{\det\Sigma}(w-\langle w\rangle)+\dfrac{z-\langle z\rangle}{\Sigma^{ZZ}}\right)$
+$=-\epsilon\left[-\dfrac{\Sigma^{WW}\Sigma^{ZW}}{\det\Sigma}+\dfrac{\Sigma^{WW}\Sigma^{ZW}}{\det\Sigma}+\dfrac{\Sigma^{ZW}}{\Sigma^{ZZ}}\right]-D^{Z}\left[-\dfrac{\Sigma^{WW}}{\det\Sigma}+\dfrac{1}{\Sigma^{ZZ}}\right]+(1-\lambda)D^{Z}\dfrac{\Sigma^{ZW}}{\det\Sigma}$
+$=\dfrac{D^{Z}}{\det\Sigma}\left[\Sigma^{WW}+(1-\lambda)\Sigma^{ZW}\right],$ (60)
+where we used $\Sigma^{ZW}=-D^{Z}/\epsilon$, which follows from (59), in the last line. Note that
+$\Sigma^{WW}+(1-\lambda)\Sigma^{ZW}$ $=\dfrac{1}{2\beta\epsilon(1-\epsilon)}\left[2D^{W}\beta\epsilon+2\beta^{2}D^{Z}+(1-\lambda)(-2\beta(1-\epsilon)D^{Z})\right]$
+$=\dfrac{D^{W}}{1-\epsilon}\left[1+\beta-(1-\lambda)(1-\epsilon)\right]$
+$>0,$ (61)
+where we used $D^{Z}:=\epsilon D^{W}$ in the second line and used the fact that $0\leq\lambda\leq 1$ and $0<\epsilon<1$ in the last inequality.
+From this relation and the fact that $\det\Sigma>0$, we conclude that
+$\dot{I}^{Z\rightarrow W}_{\lambda}$ $>0\quad\text{and}\quad\dot{I}^{W\rightarrow Z}_{\lambda}<0$ (62)
+for all $\lambda\in[0,1]$.
+Figure 3: $\tau$ dependence of $C^{WZ}(\tau)$, $C^{ZW}(\tau)$, and $|C^{WZ}(\tau)|-|C^{ZW}(\tau)|$.
+The parameter values are $\epsilon=D^{W}=1/2$, $\beta=1$.
+The characteristic period is given by $T=2\pi/\sqrt{\epsilon\beta-(1-\epsilon)^{2}/2}\simeq 10$.
+Figure 4: $\tau$ dependence of $C^{WZ}(\tau)$, $C^{ZW}(\tau)$, and $|C^{WZ}(\tau)|-|C^{ZW}(\tau)|$.
+The parameter values are $\epsilon=0.2$, $D^{W}=1/2$, $\beta=1$.
+In this case, the resonance condition $\epsilon\beta-(1-\epsilon)^{2}/2>0$ is not satisfied.
+Appendix C Comparison with cross-correlation function
+Here, we provide the results of the cross-correlation function. The cross-correlation function between turbulence $w$ and zonal flow $z$ at time lag $\tau$ is defined by
+$C^{WZ}(\tau):=\dfrac{\langle(w_{\tau}-\langle w\rangle)(z_{0}-\langle z\rangle)\rangle}{\sqrt{\Sigma^{ZZ}\Sigma^{WW}}}.$ (63)
+This function is invariant under shifting of time in the steady state.
+$C^{ZW}(\tau)$ is defined in a similar manner.
+We also investigate the difference $|C^{WZ}(\tau)|-|C^{ZW}(\tau)|$, which quantifies the directional influence between turbulence and zonal flow.
+The $\tau$-dependence of $C^{WZ}(\tau)$, $C^{ZW}(\tau)$, and the difference $|C^{WZ}(\tau)|-|C^{ZW}(\tau)|$ is shown in Figs.
+3 and 4.
+In Fig.
+3, the parameter values are the same as in Fig.
+1(c) of the main text, where the resonance condition $\epsilon\beta-(1-\epsilon)^{2}/2>0$ is satisfied with the characteristic period $T=2\pi/\sqrt{\epsilon\beta-(1-\epsilon)^{2}/2}\simeq 10$.
+Note that the correlation function $C^{ZW}(\tau)$ exhibits a peak around $\tau\simeq T/4$, which is consistent with the quasi-cycles, where the fluctuations of zonal flow follow those of turbulence with a phase lag $\sim\pi/2$.
+Notably, the difference $|C^{WZ}(\tau)|-|C^{ZW}(\tau)|$ is positive for a short time $\tau\lesssim T/4$, which suggests that turbulence is driven by zonal flow during this period.
+Because the information flow also quantifies the directional influence during a short time period, this result is consistent with the positivity of the information flow $\dot{I}^{W\rightarrow Z}_{\lambda}$.
+Figure 4 shows the case where the resonance condition $\epsilon\beta-(1-\epsilon)^{2}/2>0$ is not satisfied.
+Because there are no quasi-cycles in this case, the correlation functions do not oscillate.
+Even in this case, the difference $|C^{WZ}(\tau)|-|C^{ZW}(\tau)|$ is positive for at least a short period of time, which is also consistent with the result of the information flow.